A ball rolls up in inclined plane with an initial velocity w



(A) if its radius is r. What is it initial linear velocity? After a time t the ball stops



(B) what is it angular retardation



(C) what is it angular displacement before its stops

1
Expert's answer
2022-02-07T08:44:43-0500

Solution

in this case

a) initial angular velocity "\omega" , radius r then initial linear velocity

"v=r\omega"

b) after time t

Retardation

"\alpha=\frac{\omega-\omega'}{t}\\=\frac{v-v'}{tr}"

c) angular displacement

"\theta=\frac{\omega^2-\omega'^2}{2\alpha}\\=\frac{v^2-v'^2}{2\alpha r^2}"